In 1997, a restaurant chain had a profit of $45,000. In 2005, the company had a profit of $2,605,000. Find the average rate of change of the chain's profit in dollars per year.
Answer by checkley77(12844)

(2,605,000-45000)/8=2,560,000/8=$320,000 is the average increase per year.